OpenStudy (trumpdonaldd):

Malloy solved the equation −5x − 16 = 8; his work is shown below. Identify the error and where it was made. −5x − 16 = 8 Step 1: −5x − 16 + 16 = 8 + 16 Step 2: −5x = 24 Step 3: negative five times x over five = twenty-four over five Step 4: x = twenty-four over five Step 1: He should have subtracted 16 from each side. Step 1: He should have divided both sides by −5. Step 3: He should have multiplied both sides by −5. Step 3: He should have divided both sides by −5.

OpenStudy (wanell):

this is the work i think should help a little

OpenStudy (wanell):

5x - 16 = 8 -5x - 16 + 16 = 8 + 16 (step 1...correct) -5x = 24 ( step 2...correct) -5x/5 = 24/-5 (step 3...incorrect..he didn't divide both sides by -5) x = -24/5

OpenStudy (trumpdonaldd):

THANKS